# Recognition of Unit Segment and Polyline Graphs is $\exists\mathbb{R}$-Complete

### v2

@@ 4.1 $\exists\mathbb{R}$-Hardness for Unit Segments
Since these segments form an induced cycle in the graph $G$, these segments bound a cell $E$. In this cycle, $\ell$ and its connector segment $d_{1}$ (the one attaching to the cycle between the probes) form a chord, thus their representations split $E$ into two parts $E_{1},E_{2}$. See Figure 12 for an illustration. Note that $r(\ell)$ must be oriented from the end contained in $E$ towards the other end, since the intersection of $r(d_{1})$ and $r(\ell)$ lies in $E$, but the intersection of $r(d_{2})$ and $r(\ell)$ cannot. Since we assume that $r(\ell)$ intersects $r(b)$ before $r(a)$, the intersection of $r(a)$ and $r(\ell)$ thus lies outside of $E$.
We consider the two probe segments $p,q$ of $\ell$, which correspond to the intersection with $a$. These segments are attached to the cycle between the outermost probe segments of $\ell$, and $\ell$ itself. Furthermore, $p$ and $q$ are not intersecting any segment bounding $E_{1}$ or $E_{2}$, in particular not $r(b)$. Thus the probe segments $p$ and $q$ are both completely contained in the interior of $E_{1}$ and $E_{2}$, respectively. However, $r(a)$ can intersect the interior of only one of $E_{1}$ and $E_{2}$, but not both, since $r(a)$ and $r(\ell)$ are line segments and their single intersection is assumed to lie outside of $E$. Since both $p$ and $q$ must intersect $r(a)$, we arrive at the desired contradiction. We thus conclude that $\mathcal{A}$ is stretchable, finishing the proof of Theorem 1.
Figure 12: Suppose for the purpose of contradiction that $\ell$ intersects $a$ after $b$. Then the important segment $a$ can only intersect one of the cells $E_{1}$ and $E_{2}$. However, $E_{1}$ and $E_{2}$ contain a segment that $a$ intersects in its interior. A contradiction.
@@ 4.2 $\exists\mathbb{R}$-Hardness for Polylines
We study polylines with $k$ bends and we assume $k>0$ is a fixed constant. Specifically, we will show that $k$-PolyLine Recognition is $\exists\mathbb{R}$-complete.
The proof for polylines works very similarly to the proof for unit segments. Since the family of polylines is a strict superset of the family of unit segments, most of our additional work is on the soundness of the proof. To be able to ensure soundness, our construction of the graph $G$ will make sure that the polylines realizing the pseudolines cannot contain any bends in some region encoding the pseudoline arrangement $\mathcal{A}$ in any realization of $G$. With this property, the argument for soundness (as in the proof of Theorem 1) will carry over straightforwardly.
@@
To create the enhanced arrangement, we first create the frame: We create $22k+$ 2 vertical chains of $c\cdot n^{2}$ line segments, for some $c$ sufficiently large. We call the set of segments forming the $i$-th such vertical chain $C_{i}$, for $i=1\dots,2k+2$, numbered from left to right. We then connect these vertical chains by two horizontal chains, one at the top and one at the bottom. All of the segments forming these chains are called frame segments, and they together bound $12k+$ 1 bounded cells. We call the leftmost such cell the canvas.
@@
Next, we introduce probes from the left of our pseudoline arrangement, as we did in the proof of Theorem 1. Note that every pair of twin pseudolines shares one set of probes. We attach the probes as well as the pseudolines and their twins to $C_{1}$ using connector arcs, making sure that each connector intersects a unique arc of $C_{1}$, and that none of these arcs are intersecting.
@@
At the very end, at $C_{2k+2}$, we use a connector arc to attach each pseudoline $p_{j}$ to the top element of $L_{2k+2,j}$, and its twin $p_{j}^{\prime}$ to the bottom element. This whole construction is illustrated in Figure 13.
@@
Given a line arrangement $\mathcal{L}$ combinatorially equivalent to $\mathcal{A}$, we want to show that $G$ is realizable by polylines. It is easy to see that given $\mathcal{L}$, the canvas and its contents (probes, connectors, pseudolines, and pseudoline twins) can be realized even with line segments (i.e., using 00 bends), as we argued already in the proof of Theorem 1. The only remaining difficulty is to argue that the extended pseudolines weaving through $C_{2},\ldots,C_{2k+1}$ and finally attaching to $C_{2k+2}$ can be realized using at most $k$ bends per pseudoline and twin.
Since between every two lanes $L_{i,j}$ and $L_{i,j+1}$ there is at least one polyline that is not part of any lane, we can make the distance between these lanes arbitrarily large. We thus only need to show that for a single pair $p_{j}$, $p_{j}^{\prime}$, we can weave the polylines through $L_{2,j},\ldots,L_{2k+2,j}$. The construction is illustrated in Figure 14: The two polylines exit the canvas in parallel. Then, going from left to right, we first use all bends of $p_{j}$, leaving its twin straight. After having gone through $L_{2,j},\ldots,L_{k+2,j}$, we leave $p_{j}$ straight and use the bends of the twin to go through the remaining lanes $L_{k+3,j},\ldots,L_{2k+1,j}$ and to have the right ordering to attach to $C_{2k+2}$ using a connector.
@@
In every realization $r(G)$, the polylines realizing all probes, pseudolines, and internal vertices of $C_{2},\ldots,C_{2k+1}$ must lie on the same side of the boundary curves given by the cycle bounded by $C_{1}$ and $C_{2k+2}$.
@@
The induced subgraph of this set of vertices is connected. The connectors attaching them to the outermost cycle together with the outermost segments of $C_{2},\ldots,C_{2k+1}$ thus form a set of connectors as defined in Definition 5. Thus the lemma follows from Lemma 6 immediately. ∎
We again assume without loss of generality that all these elements lie in the interior (bounded) cell bounded by the outermost cycle. We now know that for every pair of consecutive chords $C_{i},C_{i+1}$, any element $r(v)$ that intersects the cycle bounded by these chords must also intersect the interior boundary curve of that cycle. We can thus apply Lemma 10 to these cycles to enforce the ordering of crossings along them. From this, similarly as in the proof of soundness for unit segments (note that that proof did not use the fact that the line segments were unit), we wish to get that the arrangement obtained by picking either $r(p_{j})$ or $r(p_{j}^{\prime})$ for every $j\in[n]$ restricted to the canvas has the same combinatorial structure as $\mathcal{A}$. However, for this it remains to be shown that at least one of $r(p_{j})$ and $r(p_{j}^{\prime})$ must be a straight line within the canvas. To show this, we first show that the two polylines $r(p_{j})$ and $r(p_{j}^{\prime})$ must cross often:
@@
Within the cell enclosed by the interior boundary curve of the cycle enclosed by two consecutive sets $C_{i},C_{i+1}$, for $12\leq i\leq 2k+$ 1, each $r(p_{j})$ and $r(p_{j}^{\prime})$ must cross.
@@
By Lemma 10 and by construction of $G$, the geometric order of $r(p_{j})$ and $r(p_{j}^{\prime})$ along the interior boundary curve of the cell bounded by $C_{i}$ and $C_{i+1}$ must be alternating (non-nesting). Thus, the two polylines $r(p_{j})$ and $r(p_{j}^{\prime})$ must cross within this cell. ∎
@@
Let $c,d$ be two polylines that intersect in exactly $t$ points $i_{1},\ldots,i_{t}$, with both polylines visiting the intersection points in this order. Then $c$ and $d$ have at least $t-1$ bends in total between the first and the last intersection.
@@
It is easy to see that there must be at least one bend between any two consecutive intersection points. See Figure 15 for an illustration. ∎
@@
The two twin polylines must have at least $2k$ intersection points that occur in the same order along the two twin: to find these points, we pick one per cell enclosed between $C_{i},C_{i+1}$ for $12\leq i\leq 2k+$ 1 as guaranteed by Lemma 12. Thus by Lemma 13, the two polylines have at least $12k-$ 1 bends in total outside of the canvas. This implies that at least one of them is straight inside the canvas. ∎
We conclude that $\mathcal{A}$ must be stretchable, finishing the proof of Theorem 2.

### v3

@@ 4.1 $\exists\mathbb{R}$-Hardness for Unit Segments
Since these segments form an induced cycle in the graph $G$, these segments bound a cell $E$. In this cycle, $\ell$ and its connector segment $d_{1}$ (the one attaching to the cycle between the probes) form a path of length two between two vertices of the cycle. Since $d_{1}$ is a connector segment, due to the previous observation about the fixed order of connector segments, $d_{1}$ lies between the outer two connector segments. Thus $r(d_{1})\cup r(\ell)$ split $E$ into two parts $E_{1},E_{2}$. See Figure˜13 for an illustration. Note that $r(\ell)$ must be oriented from the end contained in $E$ towards the other end since the intersection of $r(d_{1})$ and $r(\ell)$ lies in $E$, but the intersection of $r(d_{2})$ and $r(\ell)$ cannot. Since we assume that $r(\ell)$ intersects $r(b)$ before $r(a)$, the intersection of $r(a)$ and $r(\ell)$ thus lies outside of $E$.
We consider the two probe segments $p$ and $q$ of $\ell$, which correspond to the intersection with $a$. These segments are attached to the cycle between the outermost probe segments of $\ell$, and $\ell$ itself. Furthermore, $p$ and $q$ are not intersecting any segment bounding $E_{1}$ or $E_{2}$, in particular not $r(b)$. Thus the probe segments $p$ and $q$ are both completely contained in the interiors of $E_{1}$ and $E_{2}$, respectively. However, $r(a)$ can intersect the interior of only one of $E_{1}$ and $E_{2}$, but not both, since $r(a)$ and $r(\ell)$ are line segments and their single intersection is assumed to lie outside of $E$. Since both $p$ and $q$ must intersect $r(a)$, we arrive at the desired contradiction. We thus conclude that $\mathcal{A}$ is stretchable, finishing the proof of Theorem˜1.
Figure 13: Suppose for the purpose of contradiction that $\ell$ intersects $a$ after $b$. Then the important segment $a$ can only intersect one of the cells $E_{1}$ and $E_{2}$. However, $E_{1}$ and $E_{2}$ contain the inner probe segments $p$ and $q$ in their respective interiors. Since $a$ must intersect $p$ and $q$, we have a contradiction.
@@ 4.2 $\exists\mathbb{R}$-Hardness for Polylines
We study polylines with $k$ bends and we assume $k>0$ is a fixed constant (note that the case $k=0$ corresponds to segments, for which the recognition problem has already been proven $\exists\mathbb{R}$-complete in [27]). Specifically, we will show that $k$-PolyLine Recognition is $\exists\mathbb{R}$-complete.
The proof for polylines works very similarly to the proof for unit segments. Since the family of polylines is a strict superset of the family of unit segments, most of our additional work is on the soundness of the proof. To be able to ensure soundness, our construction of the graph $G$ will make sure that the polylines realizing the pseudolines cannot contain any bends in some region encoding the pseudoline arrangement $\mathcal{A}$ in any realization of $G$. With this property, the argument for soundness (as in the proof of Theorem˜1) will carry over straightforwardly.
@@
The whole construction is illustrated in Figure˜14. To create the enhanced arrangement, we first create the frame: We create $2k+2$ vertical chains of $4n+1$ line segments each. We call the set of segments forming the $i$-th such vertical chain $C_{i}$, for $i=1\dots,2k+2$, numbered from left to right. We then connect these vertical chains by two horizontal chains, one at the top and one at the bottom. More precisely, the top chain connects all the top segments of $C_{i}$ for $i=1\dots,2k+2$ such that two segments have distance three. The bottom chain behaves analogously. All of the segments forming these chains are called frame segments, and they together bound $2k+1$ bounded cells. We call the leftmost such cell the canvas.
@@
Next, we introduce probes from the left of our pseudoline arrangement, as we did in the proof of Theorem˜1. Note that every pair of twin pseudolines shares one set of probes. We attach the probes as well as the pseudolines and their twins to $C_{1}$ using connector arcs, making sure that each connector intersects a unique arc of $C_{1}$, and that none of these arcs are intersecting.
@@
At the very end, at $C_{2k+2}$, we use a connector arc to attach each pseudoline $p_{j}$ to the top element of $L_{2k+2,j}$, and its twin $p_{j}^{\prime}$ to the bottom element.
@@
Given a line arrangement $\mathcal{L}$ combinatorially equivalent to $\mathcal{A}$, we want to show that $G$ is realizable by polylines. It is easy to see that given $\mathcal{L}$, the canvas and its contents (probes, connectors, pseudolines, and pseudoline twins) can be realized even with line segments (i.e., using 0 bends), as we argued already in the proof of Theorem˜1. The only remaining difficulty is to argue that the extended pseudolines weaving through $C_{2},\ldots,C_{2k+1}$ and finally attaching to $C_{2k+2}$ can be realized using at most $k$ bends per pseudoline and twin.
Since between every two lanes $L_{i,j}$ and $L_{i,j+1}$ there is at least one polyline that is not part of any lane, we can make the distance between these lanes arbitrarily large. We thus need to show only that for a single pair $p_{j}$, $p_{j}^{\prime}$, we can weave the polylines through $L_{2,j},\ldots,L_{2k+2,j}$. The construction is illustrated in Figure˜15: The two polylines exit the canvas in parallel. Then, going from left to right, we first use all bends of $p_{j}$, leaving its twin straight. After having gone through $L_{2,j},\ldots,L_{k+2,j}$, we leave $p_{j}$ straight and use the bends of the twin to go through the remaining lanes $L_{k+3,j},\ldots,L_{2k+1,j}$ and to have the right ordering to attach to $C_{2k+2}$ using a connector.
@@
In every realization $r(G)$ of $G$, the polylines realizing all probes, pseudolines, and internal vertices of $C_{2},\ldots,C_{2k+1}$ must lie in the same cell of the arrangement created by the Jordan curves corresponding to the cycle bounded by $C_{1}$, $C_{2k+2}$, the top chain, and the bottom chain.
@@
The induced subgraph of this set of vertices is connected. The connectors attaching them to the outermost cycle together with the outermost segments of $C_{2},\ldots,C_{2k+1}$ thus form a set of connectors as defined in Definition˜4. Thus the lemma follows from Lemma˜5 immediately. ∎
Let us call the cell defined by Lemma˜10 the cell $F_{\text{region}}$. We again assume without loss of generality that $F_{\text{region}}$ is not the outer cell. This is because if it is the outer cell, we can perform a circular inversion. Under such a circular inversion all Jordan curves remain Jordan curves and the transformed drawing is still a valid geometric representation. We now observe that $\operatorname{trace}(F_{\text{region}})$ has at least 5 different elements, and so Lemma˜3 is applicable, and we know that the order of elements appearing along the boundary of $F_{\text{region}}$ is almost fixed. In particular, this implies that the chains $C_{2},\dots,C_{2k+1}$ appear in order from left to right in the cell $F_{\text{region}}$, subdividing it into a total of $2k+1$ disjoint sub-regions, where the first such sub-region is the canvas. (Note that locally, some small cells with only few elements on the boundary might be formed by the Jordan curves, but these do not influence the correctness of our argument.) We furthermore conclude that the curves $r(p_{j})$ and $r(p_{j}^{\prime})$ must traverse $F_{\text{region}}$ and while doing so cross every of the $2k+1$ sub-regions. We can thus apply Lemma˜9 to the sub-regions to enforce the ordering of crossings along them. From this, similarly as in the proof of soundness for unit segments (note that that proof did not use the fact that the line segments were unit), we wish to get that the arrangement obtained by picking either $r(p_{j})$ or $r(p_{j}^{\prime})$ for every $j\in[n]$ restricted to the canvas has the same combinatorial structure as $\mathcal{A}$. However, for this it remains to show that at least one of $r(p_{j})$ and $r(p_{j}^{\prime})$ must be a straight line within the canvas. To show this, we first show that the two polylines $r(p_{j})$ and $r(p_{j}^{\prime})$ must cross often:
@@
In every sub-region defined by the cycle enclosed by two consecutive sets $C_{i},C_{i+1}$ for $2\leq i\leq 2k+1$, we have that for all $j\in\{1,\ldots,n\}$ the curves $r(p_{j})$ and $r(p_{j}^{\prime})$ must cross.
@@
By Lemma˜9 and by construction of $G$, the geometric order of $r(p_{j})$ and $r(p_{j}^{\prime})$ along the interior boundary curve of the cell bounded by $C_{i}$ and $C_{i+1}$ must be alternating (non-nesting). Thus, the two polylines $r(p_{j})$ and $r(p_{j}^{\prime})$ must cross within this cell. ∎
@@
Let $c$ and $d$ be two polylines that intersect in exactly $t$ points $i_{1},\ldots,i_{t}$, with both polylines visiting the intersection points in this order. Then $c$ and $d$ have at least t1t-1 bends in total between the first and the last intersection.
@@
It is easy to see that there must be at least one bend between any two consecutive intersection points. See Figure˜16 for an illustration. ∎
@@
The two twin polylines must have at least $2k$ intersection points that occur in the same order along the two twins: to find these points, we pick one per cell enclosed between $C_{i},C_{i+1}$ for $2\leq i\leq 2k+1$ as guaranteed by Lemma˜11. Thus by Lemma˜12, the two polylines have at least 2k12k-1 bends in total outside of the canvas. This implies that at least one of them is straight inside the canvas. ∎
We conclude that $\mathcal{A}$ must be stretchable, finishing the proof of Theorem˜2.